OpenStudy (anonymous):

Please help!....a root of x^2-5x-1=0 is A. 1-√29/2 B. 5-√17/2 c. 1+ √29/2 D. 5+ √17/2 E. 5 +√29/2

OpenStudy (jdoe0001):

uze das kuadratik phormula :)

OpenStudy (jdoe0001):

\(\bf \textit{quadratic formula}\\ {\color{blue}{ 1}}x^2{\color{red}{ -5}}x{\color{green}{ -1}}=0 \qquad \qquad x= \cfrac{ - {\color{red}{ b}} \pm \sqrt { {\color{red}{ b}}^2 -4{\color{blue}{ a}}{\color{green}{ c}}}}{2{\color{blue}{ a}}}\) see what you get

OpenStudy (jhannybean):

\[x^2-5x-1=0\]\[(x^2-5x)=1\]\[x^2-5x\color{reD}{+\frac{25}{4}} = 1\color{red}{+\frac{25}{4}}\]\[\left(x-\frac{5}{2}\right)^2=\frac{29}{4}\]\[x=\pm \frac{\sqrt{29}}{2} +\frac{5}{2}\]

OpenStudy (jhannybean):

To find the portion in red is like completing the quadratic formula on the LHS, \(c=\left(\dfrac{-5}{2}\right)^2 = \dfrac{25}{4}\)

OpenStudy (anonymous):

Thanks So much I just checked and its right!

OpenStudy (jhannybean):

Nice :)